OpenStudy (anonymous):

What is the percent of increase from 50 to 90? A. 40% B. 44% C. 56% D. 80%

OpenStudy (holly00d1248):

what you think

OpenStudy (inowalst):

What do you think?

OpenStudy (anonymous):

a

OpenStudy (inowalst):

Right.

OpenStudy (holly00d1248):

yep it A

OpenStudy (anonymous):

yaya but to whom i medal

OpenStudy (holly00d1248):

@inowalst he answered first

OpenStudy (nothingwasthesame):

you could tag one and the other tags the other one

OpenStudy (anonymous):

B

OpenStudy (anonymous):

ok thanks

OpenStudy (inowalst):

B is wrong. Subtract 50 from 90. You get 40. Thank you.

OpenStudy (anonymous):

np

OpenStudy (anonymous):

i am gonna post a another one

OpenStudy (anonymous):

100 / 90 = 1.1111 (1.111) * 50 = 55.6 100 - 55.6 = ~44%

OpenStudy (anonymous):

i agree i think b 44%

OpenStudy (anonymous):

Its not as simple as 90 - 50

OpenStudy (anonymous):

ok

OpenStudy (holly00d1248):

where you get 55.6 though

OpenStudy (anonymous):

100 / 90 = 1.1111 (1.111) * 50 = 55.6 100 - 55.6 = ~44%

OpenStudy (inowalst):

Why would you divide 100 by 90?

OpenStudy (anonymous):

Well, in this case 90 is equal to 100%

OpenStudy (holly00d1248):

Still dont get it

OpenStudy (holly00d1248):

@MathNinja_C go help her then she opened another question

OpenStudy (anonymous):

lhjkkjjjjjjjjhvgkjjjjjjhoihhhhhhhhhhhhhhhhhhhhhhhhhhicih hihhdjcih eccejpopcjcclcnjncjndcknkc

OpenStudy (anonymous):

sorry my dog

OpenStudy (anonymous):

If there are 25 questions on a test, how do you figure out how much each question is worth? 100 / 25 = 4 points 100 / 90 = 1.111 basicaly 1.111 points for each unit

OpenStudy (holly00d1248):

O makes sense now

OpenStudy (anonymous):

50 * 1.111 = 55.6%

OpenStudy (anonymous):

Thus, going from 55.6 to 90 (which is 100% in this case) means 100 -55.6 = ~44%

OpenStudy (anonymous):

Please vote me as best response

OpenStudy (anonymous):

ok